OpenStudy (anonymous):

the length of a rectangle is 4.8 cm more than 2 times the width. If the perimeter of the rectangle is 52.8 cm, what are its dimensions?

OpenStudy (anonymous):

you must come up with an expression that will help you get the answer

OpenStudy (anonymous):

ok. The hint i got was P=2l + 2w

OpenStudy (anonymous):

I already know p is 52.8

OpenStudy (anonymous):

and l=2w+4.8 because your length must be 4.8cm longer than twice your width

OpenStudy (anonymous):

do you see how i got that

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

52.8= 4w+4.8 ?

OpenStudy (anonymous):

i dont think it is necessary to multiply w by 2

OpenStudy (anonymous):

u should get the values for w and l before multiplying by 2

OpenStudy (anonymous):

ok then 52.8=2w+4.8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

w=24 so then add 4.8 and multiply by 2?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

24+4.8 =28.8 28.8 * 2 = 57.6/ so no

OpenStudy (anonymous):

i guess i should just subtract 52.8 and 24.

OpenStudy (anonymous):

because you could plug in l=2w+4.8 into p=2l+2w getting you 4w+9.6+2w=52.8

OpenStudy (anonymous):

try and see if that works

OpenStudy (anonymous):

that should get you the right answer and sorry for the confusion

OpenStudy (anonymous):

ok yea it did thanks.